OpenStudy (benchan2580):

How do I solve problems like y=3x+4z

OpenStudy (mww):

There's no explicit solutions. You can rearrange x or z to be the subject/parameter. 3x = y - 4z --> x = (y-4z)/3 4z = y - 3x --> z = (y-3x)/4

OpenStudy (benchan2580):

Ok thanks